# Metastability in the Dissipative Quantum Rabi Model

### v1

@@ Appendix B Spectral decomposition of the steady state
Figure 8: Distribution of $p_{i}$ in the spectral decomposition of the steady state. The inset panel shows the $Q$-representations of the bosonic mode in the eigenstates $|\psi_{i}\rangle$ of the steady state.
@@
where $|\psi_{i}\rangle$ denotes the $i$-th eigenstate with probability $p_{i}$. Assuming that the probabilities $p_{i}$} are ordered in descending order, their distribution is shown in Fig. 8. Two dominant eigenstates, $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, have equal probabilities. Their $Q$-representations, displayed in the inset, confirm that they correspond to the pair of parity-breaking states. For each eigenstate, the $Q$-representation of the bosonic mode is used as its characteristic signature. Representative examples are shown in the inset, revealing that each eigenstate captures certain features of the steady-state distribution. Moreover, several eigenstates share similar features, suggesting that they can be grouped into the same classes. As shown in the main text, we group them into four classes using a standard community detection algorithm in graph theory.

### v2

@@ Appendix B Spectral decomposition of the steady state
Figure 8: Distribution of $p_{i}$ in the spectral decomposition of the steady state. The inset shows the $Q$ representations of the bosonic mode in the eigenstates $|\psi_{i}\rangle$ of the steady state.
@@
where $|\psi_{i}\rangle$ denotes the $i$th eigenstate with probability $p_{i}$. Assuming that the probabilities $p_{i}$} are ordered in descending order, their distribution is shown in Fig. 8. Two dominant eigenstates $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ have equal probabilities. Their $Q$ representations, displayed in the inset, confirm that they correspond to the pair of parity-breaking states. For each eigenstate, the $Q$ representation of the bosonic mode is used as its characteristic signature. Representative examples are shown in the inset, revealing that each eigenstate captures certain features of the steady-state distribution. Moreover, several eigenstates share similar features, suggesting that they can be grouped into the same classes. As shown in the main text, we group them into four classes using a standard community detection algorithm in graph theory.
Appendix C steady state solution
The steady state of Eq. (12) is obtained by numerically solving the linear equation
$\sum_{\alpha,\alpha^{\prime}}L_{\tilde{\alpha},\tilde{\alpha}^{\prime};\alpha,\alpha^{\prime}}R_{\alpha,\alpha^{\prime}}=0,$ (32)
where $L$ and $R$ denote the Liouvillian and the steady state in the Liouville-space basis
$\rho_{s}=$ $\sum_{\alpha,\alpha^{\prime}}R_{\alpha,\alpha^{\prime}}|\alpha,\alpha^{\prime}\rangle\rangle,$ (33)
$\mathcal{L}=$ $\sum_{\alpha,\alpha^{\prime},\tilde{\alpha},\tilde{\alpha}^{\prime}}L_{\tilde{\alpha},\tilde{\alpha}^{\prime};\alpha,\alpha^{\prime}}|\tilde{\alpha},\tilde{\alpha}^{\prime}\rangle\rangle\langle\langle\alpha,\alpha^{\prime}|.$ (34)
Here, $\alpha=\{s,n\}$, where $s$ and $n$ label the spin up or down and the photon number, respectively, $R_{\alpha,\alpha^{\prime}}\equiv\langle\alpha^{\prime}|\rho|\alpha\rangle$, and $L_{\tilde{\alpha},\tilde{\alpha}^{\prime};\alpha,\alpha^{\prime}}\equiv\rm{Tr}[(|\tilde{\alpha}^{\prime}\rangle\langle\tilde{\alpha}|)\mathcal{L}(|\alpha\rangle\langle\alpha^{\prime}|)]$. In the numerical calculation, we impose a photon number cutoff $n\leq N_{\rm cut}$. Convergence is verified by varying $N_{\rm cut}$, which shows that $\rho_{s}$ saturates for $N_{\rm cut}=499$ with a residual error below $10^{-12}$.